OpenStudy (anonymous):

h(x)=6x-x^2 for x≥3 Express h−1(x) in terms of x.

OpenStudy (anonymous):

sorry thats express h inverse in terms of x

OpenStudy (anonymous):

Can't be done fuction is not bijective

OpenStudy (anonymous):

\[x^2-6x=y\]\[x^2-6x+3=y+3\]\[(x-sqrt(3))^2=y+3\]\[x-\sqrt(3)=\pm(y+3)^2\]\[x=\pm(y+3)^2+sqrt(3)\]

OpenStudy (anonymous):

Necessary and compulsory condition for inversion is that the function has to be bijective